# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 5. Examples
[…]
@@
Let $R$ be the root system of type $A_{1}$, $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
@@
In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{1}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{1},\mathbf{x})\alpha_{1},\quad\mathbf{x}\in V.$
Then we have
@@
$Y_{1}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
and we take
$y_{1}=e^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
We have
$(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2},$
and
@@
Therefore, here $\{y_{1}\}$ is a pencil generator, and we obtain a flat pencil
@@
Now we consider the flat coordinate $t_{1}=y_{1}/\sqrt{2}$ of $\eta$, then in the new coordinate the flat pencil has the form
@@
$x^{1}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}(\log(\sqrt{2}t_{1})).$
It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $t_{1}$. Hence the monodromy group of the generalized Frobenius manifold is given by
$\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathbb{Z}.$
We considered the construction of generalized Frobenius manifolds for different types of affine Weyl groups with appropriately chosen weights. In what follows we will give some representative examples, which are chosen based on the following points:
@@
    In Example 5.2, the basic generators $y^{1},\cdots,y^{\ell}$ are pencil generators. The covering space ¯$\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is trivial, thus
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$
@@
    In Example 5.3, the basic generators $y^{1},\cdots,y^{\ell}$ are also pencil generators. However, unlike the previous example, the monodromy group
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{\ell}$
    is a proper subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
@@
    In Example 5.4, the basic generators $y^{1},y^{2},y^{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y^{1},y^{2}$ and $y^{3}$, i.e.,
@@
    Here $a_{1},a_{2}$ are some constants, and we find three possible values for them.
@@
    In Example 5.5, $y^{1},y^{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
@@
    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight. This complicates the construction of pencil generators and the generalized Frobenius manifold structure.
These examples involve most issues that we may come across in the computations. Something interesting issues will also happen when we consider other types of affine Weyl groups. More details will be presented in our subsequent publications.
@@
We choose $\omega=\omega_{2}$. Let us introduce the coordinates $x_{1},x_{2}$ of $V$ by
@@
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
The invariant $\lambda$-Fourier polynomials are given by
@@
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$ (5.1)
We take
@@
The metric on $V$ that is induced from $\mathbb{R}^{3}$ is given by
@@
Then $y^{1},y^{2}$ are pencil generators, and the corresponding flat pencil is given by
@@
The metric $\eta$ has the flat coordinates
@@
We also have $e^{\flat}=-\mathrm{d}\log y_{2}$. Note that in this example, not only $y_{1},y_{2}$ are polynomials of $t^{1}$ and $t^{2}$, but also $t^{1},t^{2}$ are polynomials of $y_{1}$ and $y_{2}$. This implies that the monodromy between $t^{1}$, $t^{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. Therefore,
$\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))=\mathrm{Stab}(\omega_{2})\rtimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\rtimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
@@
$\sigma_{1}\colon$ $x_{1}\to x_{2}-x_{1},\quad x_{2}\to x_{2},$
$T_{1}\colon$ $x_{1}\to x_{1}+1,\quad x_{2}\to x_{2},$
$T_{2}\colon$ $x_{1}\to x_{1},\quad x_{2}\to x_{2}+1.$
@@
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3}.$
The fundamental weights are given by
@@
Since $\kappa=2$, we have
$\lambda=\mathrm{e}^{-4\pi ic},$
then the $\lambda$-Fourier invariant polynomials are given by
@@
$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{2}^{3}+t_{1}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
@@
×$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
We also have $e^{\flat}=-\frac{1}{2}\mathrm{d}\log y_{3}$.
@@
$E=\frac{1}{2}v^{1}\frac{\partial}{\partial v^{1}}+\frac{1}{2}v^{2}\frac{\partial}{\partial v^{2}}+\frac{1}{2}v^{3}\frac{\partial}{\partial v^{3}},\quad e=-\frac{1}{2v^{1}}\frac{\partial}{\partial v^{1}}-\frac{1}{2v^{2}}\frac{\partial}{\partial v^{2}}-\frac{1}{2v^{3}}\frac{\partial}{\partial v^{3}}.$
@@
$x_{3}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\left(\frac{1}{2}\log(2v_{3})+\frac{1}{2}\log(2v_{2})+\frac{1}{2}\log(2v_{1})\right),$
we know that the monodromy group is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
$\mathrm{Mono}(\mathcal{M}(C_{3},\omega_{3}))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{3}~{}<~{}\mathrm{Stab}(\omega_{3})\rtimes\mathbb{Z}^{3}.$ (5.3)
Here $\mathrm{Stab}(\omega_{3})$ is nontrivial since $\sigma_{1},\sigma_{2}\in\mathrm{Stab}(\omega_{3})$. This example shows that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ is not always isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
@@
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
@@
In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. Let us consider the proper generators with the form
$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=z_{3},$
here $a_{1},a_{2}$ are two constants. Then there are three possible choices of $a_{1}$ and $a_{2}$ such that $z_{1},z_{2}$ and $z_{3}$ form a set of pencil generators, they are given by
$a_{1}=6,\quad a_{2}=-12.$ (5.4)
@@
$(\eta^{ij})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},1)$ with potential
@@
Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
@@
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},2)$ with potential
@@
We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
@@
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix},$
and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$. In all cases, the monodromy groups are given by
$\mathrm{Mono}(\mathcal{M}(B_{3},\omega_{1},k))=\mathrm{Stab}(\omega_{1})\rtimes\mathbb{Z}^{3}=\langle\sigma_{2},\sigma_{3}\rangle\rtimes\mathbb{Z}^{3}\cong D_{4}\rtimes\mathbb{Z}^{3},\quad k=1,2,3.$
@@
Since $\kappa=3$, we have
$\lambda=\mathrm{e}^{-6\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
$y_{1}=$ $\,\mathrm{e}^{-6\pi ic}\left(\mathrm{e}^{2\pi i(x_{1}+3c)}+\mathrm{e}^{-2\pi i(x_{1}+3c)}+\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{1}-x_{2}-3c)}+\mathrm{e}^{-2\pi i(x_{1}-x_{2}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}$
$+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
$y_{2}=$ $\,e^{-12\pi ic}\left(\mathrm{e}^{2\pi i(x_{2}+6c)}+\mathrm{e}^{-2\pi i(x_{2}+6c)}+\mathrm{e}^{2\pi i(3x_{1}-2x_{2}-3c)}+\mathrm{e}^{-2\pi i(3x_{1}-2x_{2}-3c)}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{2}-3x_{1}-3c)}+\mathrm{e}^{-2\pi i(x_{2}-3x_{1}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}$
$\,+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
The metric (,)(\ ,\ )( , ) has the form
@@
In the $y_{1},y_{2}$ coordinates the metric does not depend at most linearly on $\lambda$, hence $y_{1}$ and $y_{2}$ are not pencil generators. Proper generators $z_{1}$ and $z_{2}$ may be introduced by
@@
$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$
@@
In this case, $\mathrm{Stab}(\omega_{1})=\langle\sigma_{1}\rangle$, where
$\sigma_{1}(x_{1},x_{2})=(-x_{1}+x_{2},x_{2}).$
However, one has
$t_{1}=\frac{z_{1}}{2\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}},\quad t_{2}=-\frac{1}{3}\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}.$
Hence the actions of $\sigma_{1}$ on $t_{1}$ and $t_{2}$ is not trivial. In fact,
[…]
Here the value of $3]{-1}$ depends on the choice of the monodromy region of $f(z)=z^{1/3}$. Therefore we have
$\mathrm{Mono}(\mathcal{M}(G_{2},\omega_{2}))=\{e\}\rtimes\mathbb{Z}^{2}.$
@@
We have $\alpha^{\vee}_{i}=\alpha_{i},i=1,2,3$, and the fundamental weights are given by
@@
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
We have the $\lambda$-invariant Fourier polynomials
@@
$y_{3}=$ $\,\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{2\pi ix_{3}}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
The metric on $V$ that is induced from $\mathbb{R}^{4}$ is given by
@@
in these coordinates the metrics $\eta$ has the form
$\left(\eta^{ij}\right)$ $=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
@@
$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{3}t_{1}^{2}-36t_{2}^{2}t_{1}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{2}t_{3}t_{1}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right.$
$\,\left.+\left(-8t_{1}^{7}-60t_{3}t_{1}^{4}+120t_{2}^{2}t_{1}^{3}-72t_{3}^{2}t_{1}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$

### v3

@@ 5. Examples
In this section, we give some typical examples for our construction. We assume that $e_{1},\dots,e_{n}$ give an orthonormal basis of the Euclidean space $\mathbb{R}^{n}$ for $n\in\mathbb{N}$ with inner product (,)($\cdot\,$$\cdot$), and the $\ell$-dimensional space $V$ will be $\mathbb{R}^{\ell}$ or a subspace of $\mathbb{R}^{n}$ for a certain $n>\ell$ endowed with the induced inner product (,)($\cdot\,$$\cdot$). For convenience of the presentation, we will use $x_{i},y_{i},z_{i},t_{\alpha}$ to denote $x^{i},y^{i},z^{i}$ and $t^{\alpha}$ respectively. We will also use $E$ to denote the Euler vector $\tilde{E}$ of the generalized Frobenius manifolds given in (3.53).
@@
Let $R$ be the root system of type $A_{1}$, and $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
@@
In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence $\lambda=\mathrm{e}^{-2\pi ic}$. The Weyl group $W(R)$ acting on $V$ is generated by
@@
$Y_{1}=\mathrm{e}^{2\pi i\left(\omega_{1},\mathbf{x}\right)}+\mathrm{e}^{2\pi i\left(\omega_{1},\sigma_{1}(\mathbf{x})\right)}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
hence
$y_{1}=\mathrm{e}^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
The metric (,)($\cdot\,$$\cdot$) on $V$ is given by $(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2}$, hence
@@
Therefore, $y_{1}$ is a pencil generator, and we obtain a flat pencil
@@
The metric $\eta$ has the flat coordinate $t_{1}=y_{1}/\sqrt{2}$. In this new coordinate the flat pencil has the form
@@
$x^{1}\big|_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}\log(\sqrt{2}\,t_{1}).$
It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $y_{1}$, and equivalently $t_{1}$, hence the monodromy groups of $\mathcal{M}_{D}$ and $\widehat{\mathcal{M}}_{D}$ are given by
$\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathbb{Z}.$
Indeed, $\mathrm{Stab}_{W}(\omega_{1})$ is trivial and we have $\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathrm{Stab}_{W}(\omega_{1})\ltimes\mathbb{Z}$.
We have applied the above-mentioned construction of generalized Frobenius manifolds to different types of affine Weyl groups with appropriately chosen weights. In what follows we will give some representative examples, which are chosen based on the following points:
@@
    In Example 5.2, the basic generators $y_{1},y_{2}$ are pencil generators, and
    $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathrm{Mono}(\mathcal{M}_{D})~\cong~\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}.$
@@
    In Example 5.3, the basic generators $y_{1},y_{2},y_{3}$ are also pencil generators, while $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is a proper subgroup of $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$.
@@
    In Example 5.4, the basic generators $y_{1},y_{2},y_{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y_{1},y_{2}$ and $y_{3}$, i.e.,
@@
    Here $a_{1},a_{2}$ are some constants, and there are three possible values for them.
@@
    In Example 5.5, $y_{1},y_{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
@@
    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight. This complicates the construction of pencil generators.
  6. 6.
    In Example 5.7, we present a case with much more complicated construction of pencil generators. Compared with the former ones, the resulting prepotential in this case is not a Laurant polynomial in flat coordinates.
More examples will be presented in our subsequent publications [20].
@@
Choose $\omega=\omega_{2}$, and introduce the coordinates $x_{1},x_{2}$ of $V$ by
@@
We have the invariant Fourier polynomials
@@
Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$ and the invariant $\lambda$-Fourier polynomials
@@
The metric on $V$ is given by
@@
Then $\{y_{1},y_{2}\}$ is a set of pencil generators, and the corresponding flat pencil is given by
@@
The metric $\eta$ has flat coordinates
@@
We also have $e=-\mathrm{grad}_{\eta}\log y_{2}$. Note that in this example, not only are $y_{1},y_{2}$ polynomials of $t_{1}$ and $t_{2}$, but $t_{1}$ and $t_{2}$ are also polynomials of $y_{1}$ and $y_{2}$. This implies that the monodromy between $t_{1}$, $t_{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., $\widehat{\mathcal{M}}_{D}=\mathcal{M}_{D}$. Therefore,
$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})=\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Stab}_{W}(\omega_{2})\ltimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\ltimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\ltimes\mathbb{Z}^{2}.$
@@
$\sigma_{1}\colon$ $x_{1}\mapsto x_{2}-x_{1},\quad x_{2}\mapsto x_{2},$
$T_{1}\colon$ $x_{1}\mapsto x_{1}+1,\quad x_{2}\mapsto x_{2},$
$T_{2}\colon$ $x_{1}\mapsto x_{1},\quad x_{2}\mapsto x_{2}+1.$
@@
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3},$
and the fundamental weights
@@
Since $\kappa=2$, we have $\lambda=\mathrm{e}^{-4\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
@@
$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{1}^{3}+t_{2}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
@@
$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
We also have $e=-\frac{1}{2}\mathrm{grad}_{\eta}\log y_{3}$.
@@
$E=\frac{1}{2}v_{1}\frac{\partial}{\partial v_{1}}+\frac{1}{2}v_{2}\frac{\partial}{\partial v_{2}}+\frac{1}{2}v_{3}\frac{\partial}{\partial v_{3}},\quad e=-\frac{1}{2v_{1}}\frac{\partial}{\partial v_{1}}-\frac{1}{2v_{2}}\frac{\partial}{\partial v_{2}}-\frac{1}{2v_{3}}\frac{\partial}{\partial v_{3}}.$
@@
$x_{3}\big|_{\lambda=0}=$ $\frac{1}{4\pi i}\left(\log(2v_{3})+\log(2v_{2})+\log(2v_{1})\right),$
we know that the monodromy group of $\widehat{\mathcal{M}}_{D}$ is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
$\mathrm{Mono}(\widehat{\mathcal{M}}_{D})~\cong~\{e\}\ltimes\mathbb{Z}^{3}~<~\mathrm{Stab}_{W}(\omega_{3})\ltimes\mathbb{Z}^{3}.$ (5.2)
Here $\mathrm{Stab}_{W}(\omega_{3})=\langle\sigma_{1},\sigma_{2}\rangle\cong S_{3}$ is not trivial. In fact, in this case we have $\widehat{\mathcal{M}}_{D}/S_{3}\cong\mathcal{M}_{D}$. This example shows that $\mathrm{Mono}(\widehat{\mathcal{M}}_{D})$ is not isomorphic to $\mathrm{Mono}(\mathcal{M}_{D})$ in general.
@@
Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
@@
In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. A set of proper generators should have the form
$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=y_{3},$
where $a_{1},a_{2}$ are two constants. We have the following three choices for $a_{1}$ and $a_{2}$:
$a_{1}=6,\quad a_{2}=-12,$ (5.3)
@@
$(\eta^{\alpha\beta})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}$.
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{1}$ with potential
@@
Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e=-\mathrm{grad}_{\eta}\log z_{1}$.
@@
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{2}$ with potential
@@
We also have $e=-\mathrm{grad}_{\eta}\log z_{1}$.
@@
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$,
and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$, hence the third generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1})_{3}$ is isomorphic to the second one.
@@
Since $\kappa=3$, we have $\lambda=\mathrm{e}^{-6\pi ic}$. The invariant $\lambda$-Fourier polynomials are given by
$y_{1}=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
$y_{2}=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
The metric on $V$ has the form
@@
In the $y_{1},y_{2}$ coordinates the metric does not depend at most linearly on $\lambda$, hence $y_{1}$ and $y_{2}$ are not pencil generators. Proper generators $z_{1}$ and $z_{2}$ can be introduced by
@@
$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$.
@@
We also have $e=-\frac{1}{3}\mathrm{grad}_{\eta}\log z_{2}$.
@@
We have $\alpha^{\vee}_{i}=\alpha_{i},~i=1,2,3$, and the fundamental weights are given by
@@
Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$. We have the invariant $\lambda$-Fourier polynomials
@@
$y_{3}=$ $\,\mathrm{e}^{2\pi ix_{3}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
The metric on $V$ is given by
@@
in these coordinates the metric $\eta$ has the form
@@
$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{1}^{2}t_{3}-36t_{1}t_{2}^{2}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{1}t_{2}t_{3}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right$.
$\,\left.+\left(-8t_{1}^{7}-60t_{1}^{4}t_{3}+120t_{1}^{3}t_{2}^{2}-72t_{1}t_{3}^{2}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$